# The Very High Energy Afterglow of Structured Jets: GW 170817 and Prospects for Future Detections

### v1

@@ Appendix C Posterior distributions of the three fits of the afterglow of GW 170817
Fig. 16 gives the posterior joint and marginalized distributions of the model parameters for the three fits of the multi-wavelength afterglow of GW 170817 described in Section 4. The three fits use the same observational constraints and the same priors on the free parameters, and differ only by the radiative treatment: synchrotron only (”no SSC”), synchrotron and SSC assuming that all IC scatterings occur in Thomson regime (”SSC (Thomson)”), and self-consistent calculation of synchrotron and SSC processes taking into account the KN attenuation (”SSC (with KN)”). It appears clearly that the ”no SSC” and ”SSC (with KN)” fits lead to similar results, whereas the parameters inferred by the ”SSC (Thomson)” fit are slighltly different (see also Table 5). This especially true for the viewing and core jet opening angles, but also for the kinetic energy and the external density. Note the correlations between these quantities. This difference is due to the fact that the SSC component in the afterglow of GW 170817 is very weak due to the KN attenuation, and is therefore overestimated in the ”SSC (Thomson)” case: see discussion in § 4.3.

### v2

@@ Appendix C Posterior distributions of the three fits of the afterglow of GW 170817
Figure 16: Posterior joint and marginalised distributions of the free model parameters, $E_{\mathrm{0,iso}}^{\mathrm{c}}$, $\theta_{\mathrm{c}}$, $\theta_{\mathrm{v}}$, $n_{\mathrm{ext}}$, $\epsilon_{\mathrm{B}}$, $\epsilon_{\mathrm{e}}$, $\zeta$, $p$ $a$, $\Gamma_{0}^{\mathrm{c}}$, $b$ for the three fits of the afterglow of GW 170817 presented in Sect. 4, which differ only by the treatment of the radiative processes. In blue, no SSC scatterings are taken into account (”no SSC”); in green, SSC scatterings are assumed to be only in the Thomson regime (”SSC (Thomson)”); in orange, the SSC scattering is depleted at high energy in the KN regime (”SSC (with KN)”). The coloured contours correspond to the $3\sigma$ confidence intervals for each model and each parameter.
Fig. 16 gives the posterior joint and marginalised distributions of the model parameters for the three fits of the multi-wavelength afterglow of GW 170817 described in Sect. 4. The three fits use the same observational constraints and the same priors on the free parameters, and differ only by the radiative treatment: synchrotron only (”no SSC”), synchrotron and SSC assuming that all IC scatterings occur in Thomson regime (”SSC (Thomson)”), and self-consistent calculation of synchrotron and SSC processes taking into account the KN attenuation (”SSC (with KN)”). It appears clearly that the ”no SSC” and ”SSC (with KN)” fits lead to similar results, whereas the parameters inferred by the ”SSC (Thomson)” fit are slighltly different (see also Table 5). This especially true for the viewing and core jet opening angles, but also for the kinetic energy and the external density. Note the correlations between these quantities. This difference is due to the fact that the SSC component in the afterglow of GW 170817 is very weak due to the KN attenuation, and is therefore overestimated in the ”SSC (Thomson)” case: see discussion in Sect. 4.3.